OpenStudy (micahm):

Refer to the triangle below. If a = 3 and b = 3, Find c to the nearest hundredth. Show all work.

OpenStudy (micahm):



OpenStudy (anonymous):

Figure it out, this is pre- school math

OpenStudy (micahm):

it's 11 grade

OpenStudy (anonymous):

A squared + B squared = C squared

OpenStudy (anonymous):

You can use this equation \( a^2 + b^2 = c^2\) \( 3^2 + 3^2 = c^2\)

OpenStudy (micahm):

then how do i solve this Find c to the nearest hundredth. Show all work.

OpenStudy (anonymous):

scheiße leicht

OpenStudy (sleepyjess):

@MDoodler gave you the equation. All you have to do is solve. \(\sf 3^2 +3^2=c^2\)

OpenStudy (micahm):

thanks

OpenStudy (sleepyjess):

Let me know what you get

OpenStudy (anonymous):

\( a^2 + b^2 = c^2\) \( 3^2 + 3^2 = c^2\) \( 18 = c^2\)

OpenStudy (micahm):

c=\[c=_+3\sqrt{2}\]

OpenStudy (anonymous):

Now for \( 18 = c^2\) take the square root of both sides to get ride of \( c^2\) and finish the problem

OpenStudy (anonymous):

correct micahm

OpenStudy (micahm):

so that's all i had to do

OpenStudy (anonymous):

Now you have to proximate to nearest hundredths

OpenStudy (micahm):

how do i do that

OpenStudy (micahm):

how do i do the next part

OpenStudy (sleepyjess):

Take the square root of both sides

OpenStudy (micahm):

which two sides

OpenStudy (sleepyjess):

both sides of the equation

OpenStudy (micahm):

of this c=+32√

OpenStudy (sleepyjess):

Your equation right now is \(\sf 18=c^2\)

OpenStudy (anonymous):

@sleepyjess He is asking how to turn Find c to the nearest hundredth. Show all work. She has not yet found to the nearest hundredth. That is what she was asking about. Sorry I am at work so I am in an out and I don't know if the person knows how to turn a square root into a decimal or if they can just use there calculator

OpenStudy (micahm):

i not good at math but if you break it down i can solve it better

OpenStudy (anonymous):

He either needs to be show how to turn a square root to a deciaml or they might be doing it this way. Sorry don't have time to explain right now but the video might help https://www.khanacademy.org/math/pre-algebra/exponents-radicals/radical-radicals/v/approximating-square-roots

OpenStudy (anonymous):

Watch the video and then times 3 times \(\sqrt(2) \) which will give you 4.2426 now round to nearest hundredths 4.24 is the answer for c